Mrs.Santos invests P21,000 at 8% interest rate.How much more must he invest at 9 1/2% interest to have an annual income of 9% of his total investments?
Answer by richwmiller(17219)

0.08*21000+0.095(x)=0.09(21000+x)
Multiply out.
1680+0.095(x)=1890+0.09x
combine like terms
0.005x=210
isolate x
x=210/0.005
Divide
x=42000.0
check
0.08*21000+0.095(x)=0.09(21000+x)
1680+0.095*42000.0=0.09(21000+42000.0)
1680+3990.0=0.09(63000.0)
5670.0=5670.0
ok